4. The sum of James’ and David’s ages is 34 years. Five years ago, the sum of twice James’ age and three times David's age was 86 years. Using an appropriate variable for each age, form a pair of simultaneous equations and solve them to find the respective ages of the two boys.

Answer by ankor@dixie-net.com(22740)

4. The sum of James’ and David’s ages is 34 years. Five years ago, the sum of twice James’ age and three times David's age was 86 years.
Using an appropriate variable for each age, form a pair of simultaneous equations and solve them to find the respective ages of the two boys.
:
let j = James' present age
let d = David's
:
"The sum of James’ and David’s ages is 34 years."
j + d = 34
" Five years ago, the sum of twice James’ age and three times David's age was 86 years."
2(j-5} + 3(d-5) = 86
2j - 10 + 3d - 15 = 86
2j + 3d - 25 = 86
2j + 3d = 86 + 25
2j + 3d = 111
using these two equations, multiply the 1st equation by 3, subtract the 2nd
3j + 3d = 102
2j + 3j = 111
----------------subtraction eliminates d
j + 0 = -9, obviously, something is wrong with this problem